Find the general solution of the DE 3y( x2 - 1) dx  (x3  8y – 3x) dy = 0


1
Expert's answer
2022-05-17T15:18:36-0400

3y(x2-1)dx+(x3+8y-3x)dy=0

M(x,y)=3y(x2-1), N(x,y)=x3+8y-3x

dM/dy=3x2-3, dN/dx=3x2-3

dM/dy=dN/dx

dF(x,y)=0

F(x,y)=C

F(x,y)="\intop" 3y(x2-1)dx+g(y)=3yx3/3-3xy+g(y)=yx3-3xy+g(y)

F'y=N(x,y)

x3-3x+g'(y)=x3+8y-3x

g'(y)=8y

g(y)="\int" 8ydy=4y2+C

Answer: x3y-3xy+4y2=C